Can you please show me the steps for solving?
x+3(x-1)=x+18

Answer by checkley77(12569)

x+3(x-1)=x+18 MULTIPLY 3(X-1)
X+3X-3=X+18 COMBINE THE X TERMS
4X-X=18+3 ADD LIKE TERMS
3X=21 DIVIDE BY 3
X=21/3 REDUCE THE FRACTION
X=7 ANS.
PROOF:
7+3(7-1)=7+18
7+3*6=25
7+18=25
25=25